OpenStudy (anonymous):

Use the difference-of-squares pattern to factor the following. 4a^2 − (1b + 3)^2

Directrix (directrix):

Difference of Squares Pattern -------------------------- (x^2 -y^2) = (x + y)(x - y) (2a)^2 - (b + 3)^2 = [(2a) + (b + 3)] [(2a) - (b +3)] = Clean those up and post results here. I'll check.

OpenStudy (anonymous):

(2a+2a)(b+3)

Directrix (directrix):

[(2a) + (b + 3)] [(2a) - (b +3)] = [2a + b + 3] [ 2a - b - 3]

OpenStudy (anonymous):

[(2a+b+3)(2a-b-3)] Can you write it like that?

Directrix (directrix):

Yes.